# Large Language Models for Autonomous Driving (LLM4AD): Concept, Benchmark, Experiments, and Challenges

### v3

@@ VI-C Security and Privacy
A fundamental security concern lies in adversarial attacks targeting the LLM’s language understanding capabilities [goodfellow2014explaining, carlini2017towards]. Unlike traditional autonomous systems where inputs are structured and validated, LLMs must process natural language commands that could contain malicious intent. For instance, carefully crafted commands could potentially trigger the model’s semantic understanding to trigger unsafe driving behaviors or override safety protocols. These attacks are particularly challenging to defend against because they operate at the semantic level, where distinguishing between approprate and malicious commands becomes highly complex.
@@
VI-D Safety
@@
When LLMs are integrated into autonomous driving systems, safety is always the first priority. In our proposed LLM4AD concept, the LLMs themselves do not directly handle safety requirements because of the potential for hallucinations in LLM outputs and the reaction time required. Hallucinations could be dangerous in safety-critical situations, and LLMs may not react quickly enough to sudden events. In our solutions, we employ a rule-based safety guarantee mechanism where LLM outputs must fall within predefined safe ranges - if outputs exceed these bounds or do not match the output format, execution is automatically blocked.
@@
However, the challenge is that ensuring system safety and reliability requires extensive validation across a wide range of operational conditions. This includes testing system responses to edge cases, verifying behavior in degraded operating conditions [he2010single, zhu2024mwformer, 10310160, li2024light], and validating the effectiveness of fallback mechanisms. The challenge extends to maintaining clear audit trails of system decisions and ensuring alignment with different regulatory requirements. This is particularly critical for systems implementing LLM-based systems, where behavior must consistently align with defined ethical guidelines while maintaining safety guarantees.
VI-E Trust and Transparency
@@
A critical challenge in integrating LLMs into autonomous driving systems lies in establishing genuine trust and transparency between human users and the autonomous vehicle. While LLMs can generate natural language explanations for vehicle decisions [marcu2024lingoqavisualquestionanswering], several fundamental challenges emerge:
@@
First, there is the risk of explanation mismatch - the LLM might generate plausible-sounding explanations that do not actually match the true decision-making process of the underlying autonomous system. This could create a false sense of transparency and potentially dangerous misconceptions about the vehicle’s capabilities and limitations. When the LLM generates both executable policies and natural language reasoning, we must ensure perfect alignment between the two to avoid misleading explanations.
@@ VI-F Personalization
VII Conclusions
In this paper, we presented a comprehensive exploration of LLM4AD, the application of Large Language Models (LLMs) for autonomous driving. Our research highlighted the significant potential of LLMs to enhance various aspects of autonomous vehicle technology, from perception and scene understanding to language interaction and decision-making. By leveraging the natural language understanding and reasoning abilities of LLMs, we demonstrated their capacity to enhance the safety, efficiency, and user experience of autonomous vehicles. Moreover, we introduced a comprehensive benchmark specifically designed to evaluate the instruction-following abilities of LLMs within the autonomous driving domain. Our experiments, conducted on real-world vehicle platforms, showed promising results, validating the effectiveness of our proposed LLM-based approaches. These findings underscored the significant impact that LLMs can have on the future of autonomous driving technology, making it safer, smarter, and more accessible for all.
[…]
@@ VII Conclusions
[Uncaptioned image] Yupeng Zhou received his B.S. degree in Mechanical Engineering from South China University of Technology in 2022. He is now an M.S. student in the College of Engineering at Purdue University, with a research focus on applied robotics, autonomous vehicles, and controls.
[Uncaptioned image] Peiran Liu received a B.S. degree in Information Systems from the University of Florida in 2023. He is currently pursuing his Ph.D. degree in the College of Engineering at Purdue University. His research interests include computer vision, human behavior predictions, and digital twins.
@@
[Uncaptioned image] Amr Abdelraouf (Member, IEEE) received his Ph.D. degree in Transportation Engineering from the University of Central Florida in 2022. He is a research scientist at Toyota Motor North America, InfoTech Labs. His research interests include applied AI/ML, advanced driver assistance systems, connected and automated vehicles, intelligent transportation systems, and digital twins.

### v4

@@ VIII-C Security and Privacy
A fundamental security concern lies in adversarial attacks targeting the LLM’s language understanding capabilities [goodfellow2014explaining, carlini2017towards]. Unlike traditional autonomous systems, where inputs are structured and validated, LLMs must process natural language commands that could contain malicious intent. For instance, carefully crafted commands could potentially trigger the model’s semantic understanding to trigger unsafe driving behaviors or override safety protocols. These attacks are particularly challenging to defend against because they operate at the semantic level, where distinguishing between appropriate and malicious commands becomes highly complex.
@@
Our real-world implementations (Sec. VI and V) mitigate some of these risks by storing user-specific interaction logs (e.g., commands $I$, policies $P$, feedback $F$) on the vehicle’s on-board ECU in a local database, rather than in the cloud. However, this does not solve the challenge of defining clear data-governance policies. Future systems must move beyond principles and implement concrete policies for: (1) Data Retention (e.g., “voice recordings are deleted immediately after transcription,” “interaction logs are deleted after 90 days”), and (2) User Control (e.g., a “privacy mode” that disables logging, and a clear interface for users to review, export, or delete their entire stored profile).
VIII-D Safety
@@
When LLMs are integrated into autonomous driving systems, safety is always the first priority. In our proposed LLM4AD concept, the LLMs themselves do not directly handle safety requirements because of the potential for hallucinations in LLM outputs and the reaction time required. Hallucinations could be dangerous in safety-critical situations, and LLMs may not react quickly enough to sudden events. In our solutions, we employ a rule-based safety guarantee mechanism where LLM outputs must fall within predefined safe ranges if outputs exceed these bounds or do not match the output format, execution is automatically blocked.
@@
However, the challenge is that ensuring system safety and reliability requires extensive validation across a wide range of operational conditions. This includes testing system responses to edge cases, verifying behavior in degraded operating conditions [he2010single, zhu2024mwformer, 10310160, li2024light], and validating the effectiveness of fallback mechanisms. This technical challenge is intrinsically linked to legal and ethical questions. For instance, if an LLM misinterprets an ambiguous user command and contributes to an accident, it creates a complex chain of liability. While our proposed architecture mitigates this by using rule-based safety checks to override or block unsafe LLM outputs, it does not solve the fundamental ethical question of responsibility. Defining the legal and ethical frameworks for accountability in these human-AI teaming scenarios remains a significant, unresolved challenge that is critical for public trust and future deployment. The challenge extends to maintaining clear audit trails of system decisions and ensuring alignment with different regulatory requirements. This is particularly critical for systems implementing LLM-based components, where behavior must consistently align with defined ethical guidelines while maintaining safety guarantees.
VIII-E Trust and Transparency
@@
A critical challenge in integrating LLMs into autonomous driving systems lies in establishing genuine trust and transparency between human users and the autonomous vehicle. While LLMs can generate natural language explanations for vehicle decisions [marcu2024lingoqavisualquestionanswering], several fundamental challenges emerge.
@@
First, there is a risk of explanation mismatch: The LLM may generate plausible-sounding explanations that do not actually match the true decision-making process of the underlying autonomous system. This could create a false sense of transparency and potentially dangerous misconceptions about the vehicle’s capabilities and limitations. When the LLM generates both executable policies and natural language reasoning, we must ensure perfect alignment between the two to avoid misleading explanations. In our proposed frameworks, the explanation (output thought $R$) is generated concurrently with the executable policy ($P$). This explanation is intended to be shown on the vehicle’s human-machine interface (HMI) display during execution, allowing the human to understand the system’s intent. However, detecting a mismatch between the plausibility of the explanation and the technical correctness of the policy remains an open challenge. Our current human-in-the-loop feedback $F$ provides a basic mechanism for users to flag such discrepancies, but an automated verification layer for this “explanation-action alignment” is a critical area for future research.
@@ VIII-F Personalization
A further challenge is the scalability of the personalization memory and the resolution of conflicting user preferences over time. A simple approach of logging all historical interactions becomes computationally unmanageable and can lead to ambiguous or contradictory guidance for the model. However, our on-board VLM framework (Section VI) is designed to address this challenge using a RAG-based memory module. This architecture avoids feeding an unmanageably large history into the model’s context. Instead, it leverages a vector database to perform a similarity search, retrieving only the top-k most relevant prior interactions based on the current command and context. This RAG-based approach directly addresses scalability by ensuring the VLM reasons over a small, contextualized subset of data, even from a history of thousands of interactions. This method also provides a mechanism for resolving conflicts; by retrieving the most similar past scenarios, the VLM can arbitrate between a few relevant (and potentially conflicting) examples to best match the user’s preference in the current specific situation .
IX Conclusions
In this paper, we presented a comprehensive framework for integrating Large Language Models into autonomous driving, referred to as LLM4AD. We introduced the overall concept and architecture, reviewed the recent advances, proposed standardized benchmarks including LaMPilot-Bench and NuPlanQA to evaluate instruction-following and multimodal reasoning abilities, and conducted extensive experiments in both simulation and real-world environments. Our real-vehicle studies demonstrated the feasibility of deploying on-cloud and on-edge LLMs for personalized decision-making and motion control, highlighting their ability to interpret human intent and adapt to diverse driving contexts.
Looking forward, we explored the emerging Vision-Language Diffusion Models (ViLaD) as a promising future trend for LLM4AD, offering new directions for reasoning, planning, and generative scene understanding. Despite significant progress, challenges remain in achieving real-time inference, safety assurance, explainability, and personalization. We believe that future development in model efficiency, multimodal reasoning, and human-autonomy teaming will enable next-generation autonomous vehicles to be more adaptive, interpretable, and human-centric.
@@ IX Conclusions
[Uncaptioned image] Sung-Yeon Park received her M.S. degree in Civil Engineering from Seoul National University in 2023. She is now pursuing a Ph.D. degree in the College of Engineering at Purdue University, with a research focus on autonomous driving for 3D/4D scene understanding and spatial reasoning.
@@
[Uncaptioned image] Yupeng Zhou (Student Member, IEEE) received his M.S. degree in Mechanical Engineering from Purdue University, West Lafayette, in 2024. He is now a Ph.D. student in the College of Engineering at Purdue University, with a research focus on applied robotics, autonomous vehicles, and controls.
@@
[Uncaptioned image] Juntong Peng (Student Member, IEEE) received his B.Eng. degree in Information Engineering from Shanghai Jiao Tong University in 2024. He is now pursuing his Ph.D. degree in the College of Engineering at Purdue University, with a research focus on autonomous driving, collaborative perception, and multi-agent intelligence.
[Uncaptioned image] Jiaru Zhang received his Ph.D. degree in Computer Science from Shanghai Jiao Tong University in 2024. He is currently a postdoctral scholar in the Institute for Physical Artificial Intelligence at Purdue University. His research interests include Bayesian deep learning, generative models, and their applications in autonomous driving.
[Uncaptioned image] Ruqi Zhang received her Ph.D. degree from Cornell University. She is currently an assistant professor in the Department of Computer Science at Purdue University. Before that, she was a postdoctoral fellow at the Institute for Foundations of Machine Learning, UT Austin. Her research focuses on machine learning, deep learning, generative modeling, probabilistic methods, and trustworthy AI.
@@
[Uncaptioned image] Amr Abdelraouf (Member, IEEE) received his Ph.D. degree in Transportation Engineering from the University of Central Florida in 2022. He was a research scientist at Toyota Motor North America, InfoTech Labs. His research interests include applied AI/ML, advanced driver assistance systems, connected and automated vehicles, intelligent transportation systems, and digital twins.